### v2

@@ Appendix B Dutch Auctions (Section 3) Appendix
We now prove that the Dutch auction is strongly shill-proof. Towards contradiction, suppose not. So, there must exist some $S$, $\xi\in S$, and $V$ such that $\left\{\overline{{V}}_{\xi}\right\}$ is selected from the menu $\mu(V,\xi)$. But by construction, this means that the auction immediately ends and the good is allocated to the shill bidder who misreported. By Lemma A.2, the revenue from this deviation is 00, which must be weakly less than any other possible transfer.
@@
Uniqueness. Towards contradiction, suppose there exists a menu rule $\tilde{\mu}\neq\mu^{D}$ that is associated with a public, strongly shill-proof, orderly, optimal auction. Therefore, there exists $V$ and $\xi$ such that $\tilde{\mu}(V,\xi)\neq\mu^{D}(V,\xi)$. Without loss, we will suppose that $V$ is the first time in the game tree that $\tilde{\mu}$ differs from $\mu^{D}$. Formally, for all $\hat{V}\supsetneq V$, $\tilde{\mu}(V,\xi)=\mu^{D}(V,\xi)$. We now proceed in cases.
Case 1 (Different Next Player Choice). Suppose $\tilde{\mu}(V,\xi)=\left\{\left(W_{\mathsf{L}},\tilde{\vec{\xi}}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\tilde{\vec{\xi}}_{\mathsf{H}}\right)\right\},$ where $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$ or $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\vec{\xi}_{\mathsf{H}}$. If $\tilde{\vec{\xi}}_{\mathsf{L}}=\vec{\xi}_{\mathsf{L}}$, then $\tilde{\vec{\xi}}_{\mathsf{H}}=\vec{\xi}_{\mathsf{H}}$ because the outcome is fully resolved once a bidder selects the high partition. So, we need only consider the case where $\tilde{\vec{\xi}}_{\mathsf{L}}\neq\vec{\xi}_{\mathsf{L}}$. By Definition 2.1, $\tilde{\vec{\xi}}_{\mathsf{H}}\neq\emptyset$ (even if the $\xi$ chooses $\left\{\overline{{V}}_{P}\right\}$). WLOG, we can assume that bidder $b_{1}$ is called first, followed by bidder $b_{2}$ with higher priority than $b_{1}$, (potentially) followed by the remaining bidders. There must exist such a $b_{2}$ because otherwise the auction calls players in the same order as the Dutch auction, which we assumed was not the case. Let $R=B\backslash\left\{b_{1}\right\}$, and for some $m$, $v_{b_{2}}=v^{m}$ and $v_{i}=v^{m-2}$ for all bidders $i\notin\left\{b_{1},b_{2}\right\}$. Taking the expression from Lemma B.1 and dividing both sides by $X_{b_{2}}$, we get
$\frac{T_{b_{2}}(v^{m};V)}{X_{b_{2}}(v^{m};V)}=v^{m}-\sum_{k:v^{j_{k}}<v^{m}}\frac{X_{b_{2}}\left(v^{j_{k}};V\right)}{X_{b_{2}}\left(v^{m};V\right)}\cdot(v^{j_{k+1}}-v^{j_{k}})<v^{m}.$
@@
(Note that there must be at least one such $k$ in the summation because otherwise $V_{b_{2}}=\left\{v^{m}\right\}$ and $b_{2}$ would not take an action.)
@@
The last choice $b_{2}$ makes is to select a partition $W$ such that $\underline{W}\geq v^{m-2}$. We can therefore apply Lemma B.2 to conclude that the transfer if $b_{1}$ reports 00 must be $\frac{T_{b_{2}}(\underline{W};V)}{X_{b_{2}}(\underline{W};V)}\leq\frac{T_{b_{2}}(v^{m};V)}{X_{b_{2}}(v^{m};V)}<v^{m}$. If bidder $b_{1}$ instead reports $v^{m}$, then bidder $b_{2}$ will win and the revenue will be $v^{m}$ and so the auction will not be strongly shill-proof. This argument also applies if only the first $n<N$ bidders are chosen in order because of the orderliness assumption.
@@
Case 2 (Different Partitions). Suppose there exists $(W,\cdot)\in\tilde{\mu}(V,\xi)$ such that $W\notin\left\{W_{\mathsf{L}},W_{\mathsf{H}}\right\}$. There are now three sub-cases: $\underline{W}\in[\rho^{*},\overline{{V}}_{-\xi})$, $\underline{W}=\overline{{V}}_{-\xi}$, and $\underline{W}\in(0,\rho^{*})$. We need not consider the case where $\underline{W}=0$ because in that case either we can consider some other choice $W^{\prime}\notin\left\{W_{\mathsf{L}},W_{\mathsf{H}}\right\}$ or $W=V_{\xi}$ which would violate Lemma A.5. We need not consider $\underline{W}>\overline{{V}}_{-\xi}$ because $V$ is the first time $\tilde{\mu}$ differs from $\mu^{D}$ and for all $V$, $\xi$, and $(\tilde{W},\cdot)\in\mu^{D}(V,\xi)$, it is the case that $\tilde{\underline{W}}\leq\overline{{V}}_{-\xi}$.
@@
Case 2a ($\underline{W}\in[\rho^{*},\overline{{V}}_{-\xi})$). In this sub-case, there exists $m^{*}$ such that $\rho^{*}\leq\underline{W}_{\ell}\leq v^{m^{*}}<\overline{{V}}_{-\xi}$. Since $V$ is the first time that $\tilde{\mu}$ differs from $\mu^{D}$, we can suppose there exists $i$ is such that $(v^{m^{*}},\xi)\rhd(\rho^{*},i)$ because otherwise the outcome would already be resolved or the player rotation would be the only difference (Case 1). Then, suppose bidder $i$ is such that $i\in R$ and $v_{i}\geq v^{m^{*}+1}$. Take bidder $\xi\in S$ to shill $v^{m^{*}}$; and for $k\notin\left\{i,\xi\right\}$, take $v_{k}=\underline{{V}}_{k}<v^{m^{*}}$. Therefore, by Lemma B.1, observe that for the last action $i$ takes, her ex-interim transfer must be higher when shill $\xi$ reports $v^{m^{*}}$ than when she reports 00. Thus by Lemma B.2, when $v^{m^{*}}<\overline{{V}}_{-\xi}$, there exists a valuation vector $v$ such that a shill bidder would want to deviate away from reporting 00—and therefore such an auction is not strongly shill-proof.
@@
Case 2b ($\underline{W}=\overline{{V}}_{-\xi}$). In this sub-case, we know $\rho^{*}\leq\underline{W}_{\ell}=\overline{{V}}_{-\xi}$. Since $V$ has been generated via a Dutch auction so far, $\xi$ is such that for all $j\neq\xi$, $(\overline{{V}}_{j},j)\rhd(\overline{{V}}_{j},\xi)$, i.e., the current player has the lowest tie-breaking priority. Letting $j\in R$ and $v_{j}=\overline{{V}}_{-\xi}$, take bidder $\xi\in S$ to report $\overline{{V}}_{-\xi}$; and for all $k\notin\left\{j,\xi\right\}$, take $v_{k}=\underline{{V}}_{k}<\overline{{V}}_{-\xi}$. As noted, $(\overline{{V}}_{j},j)\rhd(\overline{{V}}_{j},\xi)$ and so bidder $j$ is allocated the item and not shill bidder $\xi$. Therefore, by the same argument as Case 2a, shill bidding will increase revenue.
@@
Case 2c ($\underline{W}\in(0,\rho^{*})$). By Condition LABEL:*cond:rev_prin_2.LABEL:*cond:rev_prin_2_b of Lemma A.5, we can observe that $\overline{W}\geq\rho^{*}$. In particular, there exists $j$ and $v_{j},v_{j}^{\prime}\in V_{j}$ such that $(v_{j},j)\rhd(\rho^{*},\xi)$ and $(\overline{W},\xi)\rhd(v_{j}^{\prime},j)$. The second inequality is implied by Lemma 3.2 because there must be some chance that $\xi$ could win the auction in order to affect outcomes. If any bidder is ever offered a choice with $\underline{W}\geq\rho^{*}$, then the previous cases imply that a shill bidder can profitably deviate and so we only have to consider the instances where no such choices are offered. Now, suppose $R=\left\{j\right\}$. Suppose all shill bidders play the strategy of selecting the partition $\tilde{W}$ such that $0<\tilde{\underline{W}}<\rho^{*}$ if such a choice is available. Let the final move that $j$ takes to be $W^{0,j,\mathrm{last}},W^{S,j,\mathrm{last}}$ under the shill bidders’ strategy of selecting 00 and not, respectively. Similarly, define $V^{0,\mathrm{last}},V^{S,\mathrm{last}}$ as the possible values and $\tilde{t}^{0}_{j},\tilde{t}^{S}_{j}$ as the transfers under these respective strategies. Observe that it is without loss to assume that $v^{\omega}\equiv\underline{W}^{0,j,\mathrm{last}}\leq\underline{W}^{S,j,last}$ because in the latter case the shill bidders are always acting as if they have higher values than in the former case. Next, let $p_{i}^{m,c}=\mathds{P}\left[(v^{m},j)\rhd\left(v_{i},i\right)\mid V_{i}^{c,\mathrm{last}}\right]$ and define $\zeta_{i,m}=\frac{p_{i}^{m,0}p_{i}^{\omega,S}}{p_{i}^{m,S}p_{i}^{\omega,0}}$. Observe that for all $i\neq j$ and $m\leq\omega$, $\zeta_{i,m}\geq 1$ with at least one strict inequality because $\overline{{V}}_{i}^{S,\mathrm{last}}\geq\overline{{V}}_{i}^{0,\mathrm{last}}$ for all $i$ with at least one strict inequality. So, by Lemmatta B.1 and B.2, [^35] We may assume that the possible values for $j$ are sequential above the reserve otherwise we could consider $j$ as a shill bidder for some other bidder instead by the cases above.
$\tilde{t}^{S}_{j}-\tilde{t}^{0}_{j}$ $\geq\frac{T_{\xi}(v^{\omega};V^{S,\mathrm{last}})}{X_{\xi}(v^{\omega};V^{S,\mathrm{last}})}-\frac{T_{\xi}(v^{\omega};V^{0,\mathrm{last}})}{X_{\xi}(v^{\omega};V^{0,\mathrm{last}})}$
$=\sum_{k:v^{k}\in[\rho^{*},v^{\omega})}\left((v^{k+1}-v^{k})\prod_{i\neq j}\frac{p_{i}^{k,0}}{p_{i}^{\omega,0}}\right)-\sum_{k:v^{k}\in[\rho^{*},v^{\omega})}\left((v^{k+1}-v^{k})\prod_{i\neq j}\frac{p_{i}^{k,S}}{p_{i}^{\omega,S}}\right)$
@@
$=\frac{1}{\prod_{i\neq j}p_{i}^{\omega,0}p_{i}^{\omega,S}}\cdot\Bigg{(}\sum_{k:v^{k}\in[\rho^{*},v^{\omega})}(v^{k+1}-v^{k})\cdot\left(\prod_{i\neq j}p_{i}^{k,S}p_{i}^{\omega,0}(\zeta_{i,k}-1)\right)\Bigg{)}>0.$

### v3

@@ Appendix B Strongly Shill-Proof Auctions (Section 3) Appendix
We now prove that the Dutch auction is strongly shill-proof. Towards contradiction, suppose not. So, there must exist some $S$, $\xi\in S$, and $\Theta$ such that $\left\{\overline{{\Theta}}_{\xi}\right\}$ is selected from the menu $\mu(\Theta,\xi)$. But by construction, this means that the auction immediately ends and the good is allocated to the shill bidder who misreported. By Lemma B.1, the revenue from this deviation is 0, which must be weakly less than any other possible transfer.
@@
Uniqueness. Towards contradiction, suppose there exists a menu rule $\tilde{\mu}\neq\mu^{D}_{\rho^{*}}$ that is associated with a public, strongly shill-proof, orderly, optimal auction. Therefore, there exists $\Theta$ and $\xi$ such that $\tilde{\mu}(\Theta,\xi)\neq\mu^{D}_{\rho^{*}}(\Theta,\xi)$. Without loss, we will suppose that $\Theta$ is the first time in the game tree that $\tilde{\mu}$ differs from $\mu^{D}_{\rho^{*}}$. Formally, for all $\hat{\Theta}\supsetneq\Theta$, $\tilde{\mu}\left(\hat{\Theta},\xi\right)=\mu^{D}\left(\hat{\Theta},\xi\right)$. We now proceed in cases.
[…]
$\frac{T_{b_{2}}(\theta^{m};\Theta)}{X_{b_{2}}(\theta^{m};\Theta)}=\theta^{m}-\sum_{k:\theta^{j_{k}}<\theta^{m}}\frac{X_{b_{2}}\left(\theta^{j_{k}};\Theta\right)}{X_{b_{2}}\left(\theta^{m};\Theta\right)}\cdot(\theta^{j_{k+1}}-\theta^{j_{k}})<\theta^{m}.$
@@
(Note that there must be at least one such $k$ in the summation because otherwise $\Theta_{b_{2}}=\left\{\theta^{m}\right\}$ and $b_{2}$ would not take an action.)
@@
The last choice $b_{2}$ makes is to select a partition $W$ such that $\underline{W}\geq\theta^{m-2}$. We can therefore apply Lemma B.3 to conclude that the transfer if $b_{1}$ reports 0 must be $\frac{T_{b_{2}}(\underline{W};\Theta)}{X_{b_{2}}(\underline{W};\Theta)}\leq\frac{T_{b_{2}}(\theta^{m};\Theta)}{X_{b_{2}}(\theta^{m};\Theta)}<\theta^{m}$. If bidder $b_{1}$ instead reports $\theta^{m}$, then bidder $b_{2}$ will win and the revenue will be $\theta^{m}$ and so the auction will not be strongly shill-proof.
@@
Case 2 (Different Partitions). Suppose there exists $(W,\cdot)\in\tilde{\mu}(\Theta,\xi)$ such that $W\notin\left\{W_{\mathsf{L}},W_{\mathsf{H}}\right\}$. We need not consider the case where $\underline{W}=0$ because in that case either we can consider some other choice $W^{\prime}\notin\left\{W_{\mathsf{L}},W_{\mathsf{H}}\right\}$ or $W=\Theta_{\xi}$ which would violate Lemma A.4. We need not consider $\underline{W}>\overline{{\Theta}}_{-\xi}$ because $\Theta$ is the first time $\tilde{\mu}$ differs from $\mu^{D}_{\rho^{*}}$ and for all $\Theta$, $\xi$, and $(\tilde{W},\cdot)\in\mu^{D}_{\rho^{*}}(\Theta,\xi)$, it is the case that $\tilde{\underline{W}}\leq\overline{{\Theta}}_{-\xi}$. So, there are now only three sub-cases we must consider: $\underline{W}\in[\theta^{\rho^{*}},\overline{{\Theta}}_{-\xi})$, $\underline{W}=\overline{{\Theta}}_{-\xi}$, and $\underline{W}\in(0,\theta^{\rho^{*}})$.
@@
Case 2a ($\underline{W}\in[\theta^{\rho^{*}},\overline{{\Theta}}_{-\xi})$). In this sub-case, there exists $m^{*}$ such that $\theta^{\rho^{*}}\leq\underline{W}_{\ell}\leq\theta^{m^{*}}<\overline{{\Theta}}_{-\xi}$. Since $\Theta$ is the first time that $\tilde{\mu}$ differs from $\mu^{D}_{\rho^{*}}$, we can suppose there exists $i$ such that $(\theta^{m^{*}},\xi)\rhd(\theta^{\rho^{*}},i)$ because otherwise the outcome would already be resolved or the player rotation would be the only difference (Case 1). Then, suppose bidder $i$ is such that $i\in R$ and $\theta_{i}\geq\theta^{m^{*}+1}$. Take bidder $\xi\in S$ to shill $\theta^{m^{*}}$; and for $k\notin\left\{i,\xi\right\}$, take $\theta_{k}=\underline{{\Theta}}_{k}<\theta^{m^{*}}$. Therefore, by Lemma B.2, observe that for the last action $i$ takes, her ex-interim transfer must be higher when shill $\xi$ reports $\theta^{m^{*}}$ than when she reports 0. Thus by Lemma B.3, when $\theta^{m^{*}}<\overline{{\Theta}}_{-\xi}$, there exists a valuation vector $\theta$ such that a shill bidder would want to deviate away from reporting 0—and therefore such an auction is not strongly shill-proof.
@@
Case 2b ($\underline{W}=\overline{{\Theta}}_{-\xi}$). In this sub-case, we know $\theta^{\rho^{*}}\leq\underline{W}_{\ell}=\overline{{\Theta}}_{-\xi}$. Since $\Theta$ has been generated via a Dutch auction so far, the current player has the lowest tie-breaking priority, i.e., $\xi$ is such that for all $j\neq\xi$, $(\overline{{\Theta}}_{j},j)\rhd(\overline{{\Theta}}_{j},\xi)$. Letting $j\in R$ and $\theta_{j}=\overline{{\Theta}}_{-\xi}$, take bidder $\xi\in S$ to report $\overline{{\Theta}}_{-\xi}$; and for all $k\notin\left\{j,\xi\right\}$, take $\theta_{k}=\underline{{\Theta}}_{k}<\overline{{\Theta}}_{-\xi}$. As noted, $(\overline{{\Theta}}_{j},j)\rhd(\overline{{\Theta}}_{j},\xi)$ and so bidder $j$ is allocated the item and not shill bidder $\xi$. Therefore, by the same argument as Case 2a, shill bidding will increase revenue.
@@
Case 2c ($\underline{W}\in(0,\theta^{\rho^{*}})$). By Condition iib of Lemma A.4, there must be some chance that $\xi$ could win the auction in order to affect outcomes. In particular, by definition of $\tilde{x}^{*}$, there exists $j$ and $\theta_{j}\in\Theta_{j}$ such that $(\theta_{j},j)\rhd(\theta^{\rho^{*}},\xi)$ and by Lemma 3.1, there exists $\theta_{j}^{\prime}\in\Theta_{j}$ such that $(\overline{W},\xi)\rhd(\theta_{j}^{\prime},j)$. If any bidder is ever offered a choice with $\underline{W}\geq\theta^{\rho^{*}}$, then the previous cases imply that a shill bidder can profitably deviate and so we only have to consider the instances where no such choice is offered. Now, suppose $R=\left\{j\right\}$. Suppose all shill bidders play the strategy of selecting the partition $\tilde{W}$ such that $0<\tilde{\underline{W}}<\theta^{\rho^{*}}$ if such a choice is available. Let the final move that $j$ takes to be $W^{0,j,\mathrm{last}}$ and $W^{S,j,\mathrm{last}}$ under the shill bidders’ strategy of selecting 0 and not, respectively. Similarly, define $\Theta^{0,\mathrm{last}},\Theta^{S,\mathrm{last}}$ as the possible values and $\tilde{t}^{0}_{j},\tilde{t}^{S}_{j}$ as the transfers under these respective strategies. Observe that it is without loss to assume that $\theta^{\omega}\equiv\underline{W}^{0,j,\mathrm{last}}\leq\underline{W}^{S,j,last}$ because in the latter case the shill bidders are always acting as if they have higher values than in the former case. Next, for $c\in\left\{0,S\right\}$, let $p_{i}^{m,c}=\mathds{P}\left[(\theta^{m},j)\rhd\left(\theta_{i},i\right)\mid\Theta_{i}^{c,\mathrm{last}}\right]$ and define $\zeta_{i,m}=\frac{p_{i}^{m,0}p_{i}^{\omega,S}}{p_{i}^{m,S}p_{i}^{\omega,0}}$. Observe that for all $i\neq j$ and $m\leq\omega$, it is the case that $\zeta_{i,m}\geq 1$ with at least one strict inequality because $\overline{{\Theta}}_{i}^{S,\mathrm{last}}\geq\overline{{\Theta}}_{i}^{0,\mathrm{last}}$ for all $i$ with strict inequality for at least one $i$ and $m$. So, by Lemmatta B.2 and B.3, [^32] We may assume that the possible values for $j$ are sequential above the reserve otherwise we could consider $j$ as a shill bidder for some other bidder instead by the cases above.
$\tilde{t}^{S}_{j}-\tilde{t}^{0}_{j}$ $\geq\frac{T_{\xi}(\theta^{\omega};\Theta^{S,\mathrm{last}})}{X_{\xi}(\theta^{\omega};\Theta^{S,\mathrm{last}})}-\frac{T_{\xi}(\theta^{\omega};\Theta^{0,\mathrm{last}})}{X_{\xi}(\theta^{\omega};\Theta^{0,\mathrm{last}})}$
$=\sum_{\rho^{*}\leq k<\omega}\left((\theta^{k+1}-\theta^{k})\prod_{i\neq j}\frac{p_{i}^{k,0}}{p_{i}^{\omega,0}}\right)-\sum_{\rho^{*}\leq k<\omega}\left((\theta^{k+1}-\theta^{k})\prod_{i\neq j}\frac{p_{i}^{k,S}}{p_{i}^{\omega,S}}\right)$
@@
$=\frac{1}{\prod_{i\neq j}p_{i}^{\omega,0}p_{i}^{\omega,S}}\cdot\Bigg(\sum_{\rho^{*}\leq k<\omega}(\theta^{k+1}-\theta^{k})\cdot\left(\prod_{i\neq j}p_{i}^{k,S}p_{i}^{\omega,0}(\zeta_{i,k}-1)\right)\Bigg)>0.$